OpenStudy (anonymous):

how to solve 4n(square) -8n-1 = 0 using the quadratic formula

OpenStudy (anonymous):

a = 4, b = -8, and c = -1. Then, see attachment for the formula.

OpenStudy (anonymous):

but with the negative in front of the formula does that make the - 8 a positive?

OpenStudy (anonymous):

so 8- Square root of ((-8)(-8) - 4(4)(-1) = 64 - (-16) = 48

OpenStudy (anonymous):

Close. 64 - (-16) = 64 + 16. Otherwise, you were on the right track.

OpenStudy (anonymous):

80

OpenStudy (anonymous):

So 8 Square root of 80 divided by 2(4)

OpenStudy (anonymous):

(8 - sqrt(80))/8 and (8+sqrt(80))/8. Don't forget that 80 = 16*5 though, and you can factor the 16 out of the square root.

OpenStudy (anonymous):

aaaahhhhhhhhh..... (screaming) .... ok soooo why when I put it in the calc do I get -9.88854382?

OpenStudy (anonymous):

wait when i do the proper (( i got 2.118033 ... is that the answer?

OpenStudy (anonymous):

Leave it in square root form. That's if you solve for the square root.

OpenStudy (anonymous):

my brain hurts